# Mitigating Heterogeneity among Factor Tensors via Lie Group Manifolds for Tensor Decomposition Based Temporal Knowledge Graph Embedding

### v1

@@ Abstract
Jiang Li1,2 , Xiangdong Su1,2 thanks:   Corresponding Author,Yeyun Gong3, Guanglai Gao1,2 1 College of Computer Science, Inner Mongolia University, Hohhot, China 2 National & Local Joint Engineering Research Center of Intelligent Information Processing Technology for Mongolian, Hohhot, China 3 Microsoft Research Asia lijiangimu@gmail.com, {cssxd,csggl}@imu.edu.cn, yegong@microsoft.com
@@ 2 Related Work
Static Knowledge Graph Embedding Drawing inspiration from the concept of translation invariance featured in word2vec Mikolov et al. (2013), TransE assesses the relations between entities and their links by calculating the distance from $\bm{e_{s}}+\bm{e_{\hat{r}}}$ to $\bm{e_{o}}$ using standard $l_{1}$ or $l_{2}$ norms, where $\bm{e_{s}}$ and $\bm{e_{o}}$ are the vectors that represent the starting and ending entities, and $\bm{e_{\hat{r}}}$ represents the linking relation. Following TransE, TransH Wang et al. (2014), TransR Lin et al. (2015), and TransD Ji et al. (2015) introduce various mapping ways and thus refine these embeddings for better KGE representation. ComplEx Trouillon et al. (2016) employs 3-th order tensor decomposition to capture the interactions within KGs. TorusE Ebisu and Ichise (2018) utilizes a torus (a donut-shaped manifold) for its embeddings. TorusE introduces a torus, which is a compact Abelian Lie group, and defines distance functions on the torus. The torus can be considered as a collection of multiple Lie groups. Instead, we map the factor tensors to the Lie group space, thus mitigating the distributional heterogeneity among them.
Temporal Knowledge Graph Embedding In TKGE models, the temporal information is added, and the scoring function is calculated for the quadruples to assess their reasonableness. Therefore, most TKGE models use existing KGE models as a foundation. TTransE Leblay and Chekol (2018) extends TransE and encodes time stamps $\tau$ as translations same as relations. Hence, the score function of TTransE is denoted as $\phi(s,\hat{r},o,\tau)=||\bm{s}+\bm{{\hat{r}}}+\bm{\tau}-\bm{o}||_{p}$. Furthermore, TA-TransE García-Durán et al. (2018) encode timestamps based on TransE. RotateQVS Chen et al. (2022) uses quaternion embeddings to represent both entities and relations. Recently, BoxTE Messner et al. (2022) models the TKGE based on a box embedding model BoxE Abboud et al. (2020).
Tensor Decomposition Based Temporal Knowledge Graph Embedding TComplEx Lacroix et al. (2020) and TNTComplEx Lacroix et al. (2020) expand upon the ComplEx model by executing a fourth-order tensor decomposition in temporal knowledge graphs (TKGs). This method offers a more nuanced understanding of the temporal dimensions in knowledge graphs. TeLM Xu et al. (2021) utilizes the asymmetric geometric product, a method that allows for a more sophisticated and expressive representation of temporal relationships and entities. TeAST Li et al. (2023) maps relations onto Archimedean spiral timelines, and ensures that relations occurring simultaneously are placed on the same timeline, with all relations evolving over time. The above works are all based on tensor decomposition to optimize the TKGE representation. In this paper, we focus on exploring the problem of the heterogeneity of factor tensors in tensor decomposition based TKGE models.
@@ 3.2 Tensor Decomposition for TKGE
$\mathcal{Y}\sim\mathcal{X}=\sum_{r=1}^{R}\lambda_{r}\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}\otimes\mathbf{t}_{r},$ (1)
where rank $r\in\{1,...,R\}$, $\otimes$ denotes the tensor product. $\lambda_{r}$ is scaling factor. $\mathbf{u}_{r}$, $\mathbf{v}_{r}$, $\mathbf{w}_{r}$ and $\mathbf{t}_{r}$ denote the subject entity, relation, object entity and timestamp factor tensors. Tensor decomposition based TKGE methods aim to optimize the factor tensors to make $\mathcal{X}$ as close as possible to the tensor $\mathcal{Y}$ and thus achieve more accurate link prediction.
@@ 4 Methodology
$\mathcal{Y}\approx\sum_{r=1}^{R}\lambda_{r}\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}\otimes\mathbf{t}_{r},$ (2)
where $R$ is the rank of the decomposition, $\lambda_{r}$ is scaling factor, and $(\mathbf{u}_{r},\mathbf{v}_{r},\mathbf{w}_{r},\mathbf{t}_{r})\in\mathbb{R}$ are factor tensors.
@@
$\mathcal{Y}\approx\sum_{r=1}^{R}\lambda_{r}(\mathbf{B}\cdot\bm{\alpha}_{r})\otimes(\mathbf{B}\cdot\bm{\beta}_{r})\otimes(\mathbf{B}\cdot\bm{\gamma}_{r})\otimes(\mathbf{B}\cdot\bm{\delta}_{r}).$
@@
$\mathcal{Y}\approx\sum_{j=1}^{m}\lambda_{r}\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\cdot\sum_{r=1}^{R}\alpha_{jr}\beta_{jr}\gamma_{jr}\delta_{jr},$
where $\alpha_{jr}$, $\beta_{jr}$, $\gamma_{jr}$, and $\delta_{jr}$ are the scalar coefficients corresponding to the projection of the $r$-th component onto the $j$-th basis vector in their respective factor matrices. The set of basis vectors $\{\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{m}\}$ are orthogonal to each other, ensuring that each dimension represented by these vectors is independent. Both $\lambda_{r}$ and $\lambda_{j}=\sum_{r=1}^{R}\alpha_{jr}\beta_{jr}\gamma_{jr}\delta_{jr}$ can be considered as constants. Thus, we can get a reduced-rank tensor $\mathcal{Y}^{\prime}$ as
$\mathcal{Y}\approx\mathcal{Y}^{\prime}=\sum_{j=1}^{m}\lambda_{r}\lambda_{j}\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j},$ (6)
which captures the essence of the homogeneity within the factor matrices. The rank of $\mathcal{Y}^{\prime}$, denoted by $m$, is less than the original rank $R$. Thus, homogeneous factor tensors ($\mathbf{u}_{r}$, $\mathbf{v}_{r}$, $\mathbf{w}_{r}$, $\mathbf{t}_{r}$) with a low rank can effectively approximate the target quadruple tensor $\mathcal{Y}$.
@@ 7 Conclusion
In this paper, we are the first to investigate the effect of the heterogeneity among factor tensors on link prediction in tensor decomposition based TKGE models. We mainly focus on addressing the issue of heterogeneity among elements within TKGs, which is recognized as a key challenge in this domain. Moreover, similar to the majority of TKGE models, our method is unable to process new entities that are not present in the training data.
@@ Appendix A Definition and Discussion of Heterogeneity in TKG
In KGs, ‘heterogeneity’ refers to the semantic difference of the entity and relation. Similarly, in TKGs, ‘heterogeneity’ refers to the semantic difference of the entity, relation and timestamp. The heterogeneity among entities, relations, and time is as follows: (1) The heterogeneity between entities and relations is reflected in their structural roles within the graph, with entities existing as nodes and relations represented as edges between nodes. (2) The heterogeneity between entities and time is reflected in the fact that entities represent the static components of the graph, while time affects the changes in entity attributes and their relations. (3) The heterogeneity between relations and time is reflected in that relations delineate the interactions among entities, while time characterizes the temporal aspects of these interactions, specifying when they occur and their duration. Recent work Zhang et al. (2019); Li et al. (2021); Cai et al. (2022) also indicates that KGs have an intrinsic property of heterogeneity, which contains various types of entities and relations. Since TKGs extend the KG paradigm, they inherently exhibit this heterogeneity as well. Additionally, TKGs incorporate temporal information, which further contributes to time heterogeneity.

### v2

@@ Abstract
Jiang Li1,2,3 , Xiangdong Su1,2,3 thanks:   Corresponding Author, Guanglai Gao1,2,3 1 College of Computer Science, Inner Mongolia University, China 2 National & Local Joint Engineering Research Center of Intelligent Information Processing Technology for Mongolian, China 3 Inner Mongolia Key Laboratory of Multilingual Artiffcial Intelligence Technology, China lijiangimu@gmail.com, cssxd@imu.edu.cn
@@ 2 Related Work
2.1 Static Knowledge Graph Embedding
Drawing inspiration from the concept of translation invariance featured in word2vec Mikolov et al. (2013), TransE assesses the relations between entities and their links by calculating the distance from $\bm{e_{s}}+\bm{e_{\hat{r}}}$ to $\bm{e_{o}}$ using standard $l_{1}$ or $l_{2}$ norms, where $\bm{e_{s}}$ and $\bm{e_{o}}$ are the vectors that represent the starting and ending entities, and $\bm{e_{\hat{r}}}$ represents the linking relation. Following TransE, TransH Wang et al. (2014), TransR Lin et al. (2015), and TransD Ji et al. (2015) introduce various mapping ways and thus refine these embeddings for better KGE representation. ComplEx Trouillon et al. (2016) employs 3-th order tensor decomposition to capture the interactions within KGs. TorusE Ebisu and Ichise (2018) utilizes a torus (a donut-shaped manifold) for its embeddings. TorusE introduces a torus, which is a compact Abelian Lie group, and defines distance functions on the torus. The torus can be considered as a collection of multiple Lie groups. Instead, we map the factor tensors to the Lie group space, thus mitigating the distributional heterogeneity among them.
2.2 Temporal Knowledge Graph Embedding
In TKGE models, the temporal information is added, and the scoring function is calculated for the quadruples to assess their reasonableness. Therefore, most TKGE models use existing KGE models as a foundation. TTransE Leblay and Chekol (2018) extends TransE and encodes time stamps $\tau$ as translations same as relations. Hence, the score function of TTransE is denoted as $\phi(s,\hat{r},o,\tau)=||\bm{s}+\bm{{\hat{r}}}+\bm{\tau}-\bm{o}||_{p}$. Furthermore, TA-TransE García-Durán et al. (2018) encode timestamps based on TransE. RotateQVS Chen et al. (2022) uses quaternion embeddings to represent both entities and relations. Recently, BoxTE Messner et al. (2022) models the TKGE based on a box embedding model BoxE Abboud et al. (2020). TCompoundE Ying et al. (2024) employs relation-specific and time-specific compound geometric operations to enhance the modeling of temporal dynamics and relational patterns.
2.3 Tensor Decomposition Based Temporal Knowledge Graph Embedding
TComplEx Lacroix et al. (2020) and TNTComplEx Lacroix et al. (2020) expand upon the ComplEx model by executing a fourth-order tensor decomposition in temporal knowledge graphs (TKGs). This method offers a more nuanced understanding of the temporal dimensions in knowledge graphs. TeLM Xu et al. (2021) utilizes the asymmetric geometric product, a method that allows for a more sophisticated and expressive representation of temporal relationships and entities. TeAST Li et al. (2023) maps relations onto Archimedean spiral timelines, and ensures that relations occurring simultaneously are placed on the same timeline, with all relations evolving over time. The above works are all based on tensor decomposition to optimize the TKGE representation. In this paper, we focus on exploring the problem of the heterogeneity of factor tensors in tensor decomposition based TKGE models.
@@ 3.2 Tensor Decomposition for TKGE
$\mathcal{Y}\sim\mathcal{X}=\sum_{r=1}^{R}\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}\otimes\mathbf{t}_{r},$ (1)
where rank $r\in\{1,...,R\}$, $\otimes$ denotes the tensor product. $\mathbf{u}_{r}$, $\mathbf{v}_{r}$, $\mathbf{w}_{r}$ and $\mathbf{t}_{r}$ denote the subject entity, relation, object entity and timestamp factor tensors. Tensor decomposition based TKGE methods aim to optimize the factor tensors to make $\mathcal{X}$ as close as possible to the tensor $\mathcal{Y}$ and thus achieve more accurate link prediction.
@@ 4 Methodology
$\mathcal{Y}\approx\sum_{r=1}^{R}\mathbf{u}_{r}\otimes\mathbf{v}_{r}\otimes\mathbf{w}_{r}\otimes\mathbf{t}_{r},$ (2)
where $R$ is the rank of the decomposition, and $(\mathbf{u}_{r},\mathbf{v}_{r},\mathbf{w}_{r},\mathbf{t}_{r})\in\mathbb{R}$ are factor tensors.
@@
$\mathcal{Y}\approx\sum_{r=1}^{R}(\mathbf{B}\cdot\bm{\alpha}_{r})\otimes(\mathbf{B}\cdot\bm{\beta}_{r})\otimes(\mathbf{B}\cdot\bm{\gamma}_{r})\otimes(\mathbf{B}\cdot\bm{\delta}_{r}).$
@@
$\mathcal{Y}\approx\sum_{j=1}^{m}\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\cdot\sum_{r=1}^{R}\alpha_{jr}\beta_{jr}\gamma_{jr}\delta_{jr},$
where $\alpha_{jr}$, $\beta_{jr}$, $\gamma_{jr}$, and $\delta_{jr}$ are the scalar coefficients corresponding to the projection of the $r$-th component onto the $j$-th basis vector in their respective factor matrices. The set of basis vectors $\{\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{m}\}$ are orthogonal to each other, ensuring that each dimension represented by these vectors is independent. $\lambda_{j}=\sum_{r=1}^{R}\alpha_{jr}\beta_{jr}\gamma_{jr}\delta_{jr}$ can be considered as scaling constants. Thus, we can get a reduced-rank tensor $\mathcal{Y}^{\prime}$ as
$\mathcal{Y}\approx\mathcal{Y}^{\prime}=\sum_{j=1}^{m}\lambda_{j}\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j}\otimes\mathbf{b}_{j},$ (6)
which captures the essence of the homogeneity within the factor matrices. The rank $m$ of $\mathcal{Y}^{\prime}$ is less than the original rank $R$ ($m<R$). Thus, homogeneous factor tensors ($\mathbf{u}_{r}$, $\mathbf{v}_{r}$, $\mathbf{w}_{r}$, $\mathbf{t}_{r}$) with a low rank can effectively approximate the target quadruple tensor $\mathcal{Y}$.
@@ 7 Conclusion
In this paper, we investigate the effect of the heterogeneity among factor tensors on link prediction in tensor decomposition based TKGE models. We mainly focus on addressing the issue of heterogeneity among elements within TKGs, which is recognized as a key challenge in this domain. Moreover, similar to the majority of TKGE models, our method is unable to process new entities that are not present in the training data.
Acknowledgments
This work was funded by National Natural Science Foundation of China (Grant No. 62366036), National Education Science Planning Project (Grant No. BIX230343), Key R&D and Achievement Transformation Program of Inner Mongolia Autonomous Region (Grant No. 2022YFHH0077), The Central Government Fund for Promoting Local Scientific and Technological Development (Grant No. 2022ZY0198), Program for Young Talents of Science and Technology in Universities of Inner Mongolia Autonomous Region (Grant No. NJYT24033), Inner Mongolia Autonomous Region Science and Technology Planning Project (Grant No. 2023YFSH0017), Science and Technology Program of the Joint Fund of Scientific Research for the Public Hospitals of Inner Mongolia Academy of Medical Sciences (Grant No.2023GLLH0035).
@@ Appendix A Definition and Discussion of Heterogeneity in TKG
In KGs, ‘heterogeneity’ refers to the semantic difference of the entity and relation. Similarly, in TKGs, ‘heterogeneity’ refers to the semantic difference of the entity, relation and timestamp. The heterogeneity among entities, relations, and time is as follows: (1) The heterogeneity between entities and relations is reflected in their structural roles within the graph, with entities existing as nodes and relations represented as edges between nodes. (2) The heterogeneity between entities and time is reflected in the fact that entities represent the static components of the graph, while time affects the changes in entity attributes and their relations. (3) The heterogeneity between relations and time is reflected in that relations delineate the interactions among entities, while time characterizes the temporal aspects of these interactions, specifying when they occur and their duration. Recent work Zhang et al. (2019); Li et al. (2021); Cai et al. (2022) also indicates that KGs have an intrinsic property of heterogeneity, which contains various types of entities and relations. Since TKG extends the KG paradigm, they inherently exhibit this heterogeneity as well. Additionally, TKGs incorporate temporal information, which further contributes to time heterogeneity.